What Kind Of Conic Graph Is This? (x-3)^2 + (y+2)^2=4
Answer by jim_thompson5910(35256)

This is an equation of a circle since it's in the form (x-h)^2+(y-k)^2=r^2 where h = 3, k = -2 and r = 2